# ONEBench to Test Them All: Sample-Level Benchmarking Over Open-Ended Capabilities

### v1

@@ 4 ONEBench: Creation & Capability Querying
Figure 4: Constituent datasets of ONEBench-LLM (left) and OneBench-LMM (right). We provide task type, metric, and license about each dataset in Tab. 2 and Tab. 3.
After evaluating the robustness of our aggregation method across incomplete and heterogeneous measurements, we present the overall system applied to two large-scale, real-world instances of ONEBench for foundation models: LLMs and LMMs. We first outline how these benchmarks were created, then show how to use them to test arbitrary capabilities, and finally highlight key insights gained.
@@
LegalBench [43] HELM Legal 1K QEM Unknown
MATH [53] HELM Maths 1K QEM MIT
MedQA [59] HELM Medical 1K QEM MIT
NarrativeQA [68] HELM Openbook QA 1K F1 Apache-2.0
NaturalQuestions [71] HELM Search Engine Queries 1K F1 CC BY-SA 3.0
OpenbookQA [106] HELM Openbook QA 1K EM Apache-2.0
WMT 2014 [12] HELM Machine translation 1K BLEU CC-BY-SA-4.0
ARC [19] Leaderboard General QA $1.1$K EM CC-BY-SA-4.0
HellaSwag [170] Leaderboard Reasoning $10$K EM MIT
TruthfulQA [82] Leaderboard General QA $817$ EM Apache-2.0
Winogrande [129] Leaderboard Reasoning $1.2$K EM Apache-2.0
GSM8K [20] HELM + Leaderboard Maths $1.3$K QEM MIT
MMLU [52] HELM + Leaderboard General QA $13.8$K EM MIT
@@
Chatbot Arena [18] Chatbot Arena Pairwise Battles $51$K - CC BY 4.0
@@
A-OKVQA [133] VHELM VQA $7.2$K QEM Apache-2.0
Bingo [24] VHELM Bias+Hallucination $886$ ROUGE Unknown
Crossmodal-3600 [147] VHELM Captioning $1.5$K ROUGE CC BY-SA 4.0
Hateful Memes [64] VHELM Hate Speech $1$K QEM Custom(Meta)
Mementos [160] VHELM Sequential Reasoning $945$ GPT CC-BY-SA-4.0
MultipanelVQA [33] VHELM VQA $200$ QEM MIT
OODCV-VQA [152] VHELM VQA $1$K QEM CC-BY-NC-4.0
PAIRS [35] VHELM Bias $508$ QEM Unknown
Sketchy-VQA [152] VHELM VQA $1$K QEM CC-BY-NC-4.0
AI2D [61] LMMs-Eval Maths+Science $3.09$K QEM Apache-2.0
IconQA [88] LMMs-Eval Docs and Infographics $43$K ANLS CC BY-SA 4.0
InfoVQA [101] LMMs-Eval Docs and Infographics $6.1$K ANLS Unknown
LLaVA-in-the-Wild [85] LMMs-Eval Multi-disciplinary $60$ GPT4 Apache-2.0
ChartQA [99] LMMs-Eval Docs and Infographics $2.5$K QEM GPL-3.0
CMMMU [173] LMMs-Eval Multi-disciplinary $900$ QEM CC-BY-4.0
DocVQA [100] LMMs-Eval Docs and Infographics $10.5$K ANLS Unknown
MMBench [86] LMMs-Eval Multi-disciplinary $24$K GPT Apache-2.0
MMVET [167] LMMs-Eval Multi-disciplinary $218$ GPT Apache-2.0
MP-DocVQA [150] LMMs-Eval Docs and Infographics $5.2$K QEM MIT
NoCaps [3] LMMs-Eval Captioning $4.5$K ROUGE MIT
OK-VQA [98] LMMs-Eval VQA $5.1$K ANLS Unknown
RefCOCO [60, 97] LMMs-Eval Captioning $38$K ROUGE Apache-2.0
ScienceQA [89] LMMs-Eval Maths+Science $12.6$K EM CC BY-NC-SA 4.0
TextCaps [139] LMMs-Eval Captioning $3.2$K ROUGE CC BY 4.0
TextVQA [140] LMMs-Eval VQA $5$K EM CC BY 4.0
COCO [83] VHELM+LMMs-Eval Captioning $45.5$K ROUGE CC-BY-4.0
Flickr30k [165] VHELM+LMMs-Eval Captioning $31$K ROUGE CC-0 Public Domain
GQA[56] VHELM+LMMs-Eval Scene Understanding $12.6$K QEM CC-BY-4.0
MathVista [90] VHELM+LMMs-Eval Maths+Science $1$K QEM/GPT4 CC-BY-SA-4.0
MME [36] VHELM+LMMs-Eval Multi-disciplinary $2.4$K QEM/C+P Unknown
MMMU [169] VHELM+LMMs-Eval Multi-disciplinary $900$ QEM CC BY-SA 4.0
POPE [78] VHELM+LMMs-Eval Hallucination $9$K QEM/EM MIT
SEED-Bench [73, 74] VHELM+LMMs-Eval Multi-disciplinary $42.5$K QEM/EM Apache
VizWiz [44] VHELM+LMMs-Eval VQA $4.3$K QEM/EM CC BY 4.0
VQAv2 [42] VHELM+LMMs-Eval VQA $214$K QEM/EM CC BY 4.0
@@
Vision Arena [91] - Pairwise Battles $9$K - MIT
LMMs-Eval(Prometheus2) [66] - Pairwise Battles $610$K - MIT
@@
Table 3: Datasets in ONEBench-LMM: a diverse collection of benchmarks testing the abilities of LLMs in tasks such as general VQA, image captioning, hate speech detection, bias and hallucination understanding, maths and science, documents and infographics, scene understanding and sequential reasoning as well as the performance of LMMs in pairwise battles.
4.1 Creation of ONEBench-LLM & ONEBench-LMM
We develop two instantiations of ONEBench: (1) ONEBench-LLM for language models and (2) ONEBench-LMM for multimodal (vision and language) models, as outlined in  Fig. 4.
4.1.1 ONEBench-LLM
Data Pool $\mathcal{D}$. For ONEBench-LLM (Tab. 2), we source data from the Open LLM Leaderboard [8], HELM [79], and LMArena [18]. Open LLM Leaderboard and HELM aggregate several individual benchmarks (such as MMLU [52] and HellaSwag [170]), while LMArena uses pairwise model comparisons based on user-generated prompts, with user votes determining the superior model. Metrics which are converted to sample-wise ordinal rankings here include F1-Score, Exact Match (EM), and Quasi-Exact Match (QEM), as well as pairwise preferences from LMArena.
Models $\mathcal{F}$. For ONEBench-LLM, we use the 100 most downloaded models from Open LLM Leaderboard and all 79 models from HELM (as of v1.9.0), including both proprietary models like GPT-4o [112] and open-weights ones like LLaMA-3 [105]. A full list of evaluated models is provided in Appx. A.
4.1.2 ONEBench-LMM
Data Pool $\mathcal{D}$. For ONEBench-LMM (Tab. 3), data is sourced from VHELM, LMMs-Eval, and WildVisionArena. Similar to ONEBench-LLM, VHELM and LMMs-Eval aggregate individual datasets like MMMU [169] and VQAv2 [42], while WildVisionArena uses pairwise tests for LMMs through image-based chats. We convert a diverse set of metrics to sample-wise rankings, from binary metrics like EM, QEM, to real-valued scores like ROUGE [81], Perception (P), and Cognition (C) scores from MME [36]. We additionally combine pairwise comparisons from WildVisionArena with LLM-As-A-Judge preferences generated using Prometheus-2 [66], which correlate highly with human judgments. The preference comparisons are sampled randomly from LMMs-Eval while avoiding overlap with cardinal measurements.
Models $\mathcal{F}$. For ONEBench-LMM, we use 14 models from LMMs-Eval [176] and 25 models from VHELM [23], including proprietary models like Gemini Pro Vision [146] and open-weights models like LLaVA [85]. A complete list of evaluated models is provided in Appx. A.
4.2 Capabilities and Concept Probing: Results & Insights
Benchmark ##\##Concepts Cohen-$\kappa$ mAP CMC@1 CMC@10
ONEBench-LLM 40 0.793 0.8462 0.95 1.0
ONEBench-LMM 50 0.912 0.7337 0.94 0.96
Table 4: Capability Probing (Quantitative): we provide a summary of the number of concepts curated for capability probing, along with the inter-annotator agreement and retrieval metrics.
Here, we present empirical results on generating arbitrary test sets and rankings. Our goal is to enable users to make targeted queries within ONEBench and interactively search across the data pool, helping them identify the best models for their needs. To achieve this, we extend our system with a flexible mechanism for personalized aggregation, allowing users to (1) retrieve relevant data instances through structured filters and semantic search, and (2) dynamically generate model rankings based on the retrieved samples.
Setup. Given a query, the system retrieves relevant data samples using a combination of flexible semantic search and rigid constraints. This concept querying mechanism provides a personalized comparison of foundation model capabilities. We use two querying mechanisms. (i) Semantic search: we perform k-NN lookup in the embedding space of all-MiniLM-L6-v2 [123] for language tasks and SigLIP-B16 [172] for vision-language tasks, using cosine similarity. We retrieve the top-k samples for a given concept with a tuned cut-off similarity score of $0.3$ and $0.7$ for ONEBench-LLM and ONEBench-LMM, respectively. (ii) Metadata search: we verify that per-sample metadata satisfies the constraints defined in the query. Some benchmarks, such as MMMU, are equipped with detailed metadata, including categories like image type (‘diagram’, ‘table’, etc.), question type (‘multiple-choice’), field (‘Electronics’), and subfield (‘Analog Electronics’). Other benchmarks, such as image captioning datasets (e.g. COCO, NoCaps, Flickr30k), contain little metadata.
With these resources, we gather representative samples for individual queries across diverse test sets, and aggregate the ordinal sample-level model rankings using the Plackett-Luce model to produce a model ranking for that particular query.
Concepts Tested. We curated a diverse set of 50 concepts to test the breadth and versatility of ONEBench, ranging from domain-specific knowledge, such as the Coriolis Effect, to broader academic disciplines like Neuroscience, and everyday consumer goods like the Apple iPad. We showcase 6 of them in the main paper and present the rest in Appx. B. We present the results from concept querying in Figure 5 and summarize our insights below:
Insight 1. Are the retrieved data samples accurate? Two expert annotators manually reviewed and filtered out incorrect matches  [^4] The inter-annotator agreement, measured by Cohen’s Kappa, is shown in Table 4, with high values of 0.793 and 0.912 indicating strong consistency between annotators.. To evaluate the quality of the retrieved samples, we report average precision (AP) scores for a random subset of queried concepts in Fig. 5, with a full list of scores in Appx. B. Aggregating over all tested concepts in Table 4, our mAP over the concepts is 0.84 and 0.73 for ONEBench-LLM and ONEBench-LMM, respectively, demonstrating that we can reliably retrieve samples that match the intended capabilities, although there is substantial scope for improvement in some cases. Please refer to the per-concept AP in Tab. 5 for a better indicator of underrepresented concepts in ONEBench. Note that the retrieval mechanism is expected to only improve with better retrieval models, larger test sets that cover a more diverse range of capabilities, and the integration of more sophisticated querying mechanisms in ONEBench.
Insight 2. Do models perform differently across queries? A key check is to verify whether models perform distinctly across different capability queries. If the results are similar regardless of the query, fine-grained querying may be less useful, as the top model from a generic leaderboard could be a good candidate for any specific capability, as is common practice currently. However, we observe in Fig. 5 and Fig. 6 that different models perform well on different domains and concepts. This suggests that ONEBench returns valid candidate models for arbitrary user queries.
5 Related Works
Multi-task Benchmarks as Broad Capability Evaluators. Multi-task leaderboards have been the standard for benchmarking foundation models. Examples include GLUE [158], decaNLP [104], SuperGLUE [157], BigBench [142], Dynabench [65], Open LLM Leaderboard [8], CLIP-Benchmark [72], ELEVATOR [75], StableEval [153] and DataComp-38 [37], as well as massive multitask benchmarks like XTREME [138] and ExT5 [5]. However, concerns have arisen regarding the limitations of multi-task benchmarks [14]. Issues include saturation and subsequent discarding of samples [80, 11, 113, 31, 162], susceptibility to dataset selection [25], obscuring progress by evaluation metrics [22, 132], training on test tasks [154, 28, 109, 107, 143, 159], and data contamination [29, 96, 41, 26, 127, 40, 128]. ONEBench tackles these challenges by enabling extensive reuse of samples for broader model comparisons, avoiding task selection bias through democratized sourcing of samples, and using ordinal rankings to avoid evaluation minutia. Sample-level evaluation with sparse inputs also allows selective removal of contaminated data for fairer comparisons. Moreover, by supporting over-ended, evolving evaluation, it makes it harder to train on all test tasks, as opposed to fixed leaderboards that are easier to game.
On Aggregation across Benchmarks. The dominant approach to benchmarking has traditionally been multi-task benchmarks, where the most common aggregation strategy is the arithmetic mean of scores across individual tasks. However, this approach assumes that the scoring metrics are homogeneous and scaled correctly, and treat tasks of different complexities equally [108, 117]. In consequence, simple normalization preprocessing influences the rankings [21], and makes them nearly entirely dependent on outlier tasks [2]. Simply changing the aggregation method from arithmetic to geometric or harmonic mean can change the ranking [136]. Similarly, including irrelevant alternative models can change statistical significance or even change the ranking entirely [9, 174]. Mean-aggregation also has significant failure modes in handling missing scores in benchmarks [54]. The benchmarking paradigm is hence shifting towards adopting evaluation principles from other fields, such as non-parametric statistics and social choice theory [16, 125]. We use ordinal rankings instead of scores, similar to LMArena. However, unlike Arena, we use the pairwise variant of the Plackett-Luce model, which has been shown to have advantages both theoretically and empirically [116]. We benefit from some of its theoretical properties like identifiability, sample-efficient convergence, provable robustness to irrelevant alternatives, non-dominance of outliers and empirical robustness across a wide range of real-world factors which affect ranking. Moreover, we do not aggregate over benchmarks in the first place—our primary proposal is to avoid monolithic benchmarks and consider aggregation on a sample level, needing to tackle incomplete and heterogeneous measurements. We note that several other social-choice theory-based models such as score-based models [125, 137] based on the Condorcet-winner criterion [164] have been proposed, yet they were primarily applied for aggregation on multi-task benchmarks, whereas a crucial component of our proposal is to break down the benchmark boundaries and aggregate heterogeneous samples.
Dynamic Evaluation and Active Testing. Some previous works like [58, 69, 70, 130, 55, 178] tackle the ‘active testing’ problem, where the goal is to identify small “high-quality” test data-subsets, from a large pool of uncurated evaluation data. These works typically assume that the cost of unlabeled test data acquisition is low whereas the cost of acquiring per-instance labels is high. However, as pointed out in [122], these assumptions are unrealistic for foundation models, as both the acquisition of test data and label annotations can be tedious in general. Hence, in our work, we tackle a broader problem: given a large testing data pool, how can we curate and query to produce a consistent and targeted set of model rankings?
Efficient Evaluation. As evaluation suites have grown, associated inference costs have also increased. Recent research has focused on creating compressed subsets of traditional benchmarks to address this issue [155, 120, 156, 177, 115, 67, 114]. Popular approaches include subsampling benchmarks to preserve correlations with an external source like LMArena [111], sample clustering to gauge sample difficulty and then sub-sampling[156], item-response-theory based methods for informatively sampling a subset of samples for evaluation[120], or designing evolving sample-level benchmarks [122]. While the work of Prabhu et al. [122] is similar to us in principle, it requires binary metrics as input and does not handle incomplete input matrices, which is necessary for aggregation over multiple time steps. We precisely address these limitations by showing efficient evaluation while accommodating incomplete data and extending it to ordinal ranks.
Democratizing Evaluation. Standard image classification and retrieval benchmarks are collected from platforms like Flickr, which are predominantly Western-centric [4, 135]. This has raised the important question: “Progress for whom?”, with many seminal works showcasing large disparities in model performance on concepts [51, 110], tasks [47, 48, 49], and even input samples [45, 121, 144] from the Global South. In response, works have developed benchmarks tailored to diverse cultures and demographics to include their voice in measuring progress [92, 110, 121, 118]. Further works have tried to create personalized, task-specific benchmarks for flexibly evaluating models based on user-preferences [17, 77, 131, 168]— Zhang et al. [175] created Task-Me-Anything that enables users to input specific queries that then get processed to provide model rankings or responses to the query. However, their system is entirely procedurally generated, thereby not reflecting the real-world use-cases that models are typically subjected to in practice. Further, they are restricted to the fixed set of instances in their task generator pool. We take a different approach by creating flexible benchmarks where individuals, and contributing entities, can add their own samples and preferences collected from both real-world benchmarks and live model arenas like LM-Arena, thereby providing users with a realistic overview of model rankings on practical scenarios. Further, during capability testing, users can select similar preferences, making ONEBench more inclusive than traditional test sets.
6 Conclusions and Open Problems
This work tackles scalable benchmarking of arbitrary capabilities of foundation models, requiring a shift from traditional fixed training and test splits, by introducing ONEBench, an open-ended benchmarking framework for foundation models. Our open-source, democratized benchmarking methodology allows diverse evaluation samples and model measurements with detailed metadata. This affords creating customized benchmarks and testing arbitrary capabilities, including using semantic and structured searches. We provide a principled aggregation mechanism, that is both theoretically grounded and empirically validated to be robust to incomplete data and heterogeneous measurements across evaluations. We demonstrate the utility of ONEBench in two domains: ONEBench-LLM and ONEBench-LMM, showing how dynamic probing reveals new insights into model performance on specific tasks, domains, or concepts. This combination of theoretical rigour, empirical results, and practical flexibility makes ONEBench a valuable tool for comprehensively evaluating foundation models. We provide some promising directions for improvement below:
  1. 1.
  2. 2.
  3. 3.
    Structured Querying and Enhanced Retrieval: One can improve retrieval by better querying mechanisms using models like ColBERT [62] and ColPALI [34], further optimized using DSPy [63]. A particularly interesting direction is allowing compositional queries, where users combine multiple queries to test behaviour in foundation models, similar to works like ConceptMix [161] and SkillMix [166].
  4. 4.
    On the Limits of Capability Probing: While we currently allow broad, open-ended inputs to probe capabilities, some are easier to assess than others [95, 76]. As foundation models become more generalizable, a thorough analysis identifying which capabilities can be easily, reliably evaluated, which are possible to evaluate but challenging, and which are in principle impossible to evaluate is needed—this will help improve benchmarking effectiveness.
Acknowledgements
The authors would like to thank (in alphabetic order): Nikhil Chandak, Karel D’Oosterlinck, Shashwat Goel, Shyamgopal Karthik, Bálint Mucsányi and Tim Weiland for their helpful suggestions and feedback. Special thanks to Bo Li for (promptly and consistently) providing sample-level model evaluation logs of LMMs-Eval, Yujie Lu for providing early access to the battle data of Vision Arena and Palzer Lama for creating the logo of ONEBench and helping with figures. The authors would also like to thank Federico D’Agostino for recommending the title of the paper and the name of the benchmark. AP and MB acknowledge financial support via the Open Philanthropy Foundation funded by the Good Ventures Foundation. VU and SD thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) and the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VU is supported by a Google PhD Fellowship in Machine Intelligence. MB acknowledges support via the CRC 1233 on Robust Vision and from the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645.
Appendix A Models used in ONEBench:Further Details

### v2

@@ Appendix A Datasets used in ONEBench: Further Details
Figure 5: Constituent datasets of ONEBench-LLM (left) and OneBench-LMM (right). We provide task type, metric, and license about each dataset in table 5 and table 6.
@@
LegalBench (Guha et al., 2024) HELM Legal 1K QEM Unknown
MATH (Hendrycks et al., 2021b) HELM Maths 1K QEM MIT
MedQA (Jin et al., 2021) HELM Medical 1K QEM MIT
NarrativeQA (Kočiskỳ et al., 2018) HELM Openbook QA 1K F1 Apache-2.0
NaturalQuestions (Kwiatkowski et al., 2019) HELM Search Engine Queries 1K F1 CC BY-SA 3.0
OpenbookQA (Mihaylov et al., 2018) HELM Openbook QA 1K EM Apache-2.0
WMT 2014 (Bojar et al., 2014) HELM Machine translation 1K BLEU CC-BY-SA-4.0
ARC (Clark et al., 2018) Leaderboard General QA $1.1$K EM CC-BY-SA-4.0
HellaSwag (Zellers et al., 2019) Leaderboard Reasoning $10$K EM MIT
TruthfulQA (Lin et al., 2022) Leaderboard General QA $817$ EM Apache-2.0
Winogrande (Sakaguchi et al., 2021) Leaderboard Reasoning $1.2$K EM Apache-2.0
GSM8K (Cobbe et al., 2021) HELM + Leaderboard Maths $1.3$K QEM MIT
MMLU (Hendrycks et al., 2021a) HELM + Leaderboard General QA $13.8$K EM MIT
@@
Chatbot Arena Chiang et al. (2024) Chatbot Arena Pairwise Battles $51$K - CC BY 4.0
@@
A-OKVQA (Schwenk et al., 2022) VHELM VQA $7.2$K QEM Apache-2.0
Bingo (Cui et al., 2023) VHELM Bias+Hallucination $886$ ROUGE Unknown
Crossmodal-3600 (Thapliyal et al., 2022) VHELM Captioning $1.5$K ROUGE CC BY-SA 4.0
Hateful Memes (Kiela et al., 2020) VHELM Hate Speech $1$K QEM Custom(Meta)
Mementos (Wang et al., 2024b) VHELM Sequential Reasoning $945$ GPT CC-BY-SA-4.0
MultipanelVQA (Fan et al., 2024) VHELM VQA $200$ QEM MIT
OODCV-VQA (Tu et al., 2023) VHELM VQA $1$K QEM CC-BY-NC-4.0
PAIRS (Fraser and Kiritchenko, 2024) VHELM Bias $508$ QEM Unknown
Sketchy-VQA (Tu et al., 2023) VHELM VQA $1$K QEM CC-BY-NC-4.0
AI2D (Kembhavi et al., 2016) LMMs-Eval Maths+Science $3.09$K QEM Apache-2.0
IconQA (Lu et al., 2021) LMMs-Eval Docs and Infographics $43$K ANLS CC BY-SA 4.0
InfoVQA (Mathew et al., 2022) LMMs-Eval Docs and Infographics $6.1$K ANLS Unknown
LLaVA-in-the-Wild (Liu et al., 2023a) LMMs-Eval Multi-disciplinary $60$ GPT4 Apache-2.0
ChartQA (Masry et al., 2022) LMMs-Eval Docs and Infographics $2.5$K QEM GPL-3.0
CMMMU Zhang et al. (2024a) LMMs-Eval Multi-disciplinary $900$ QEM CC-BY-4.0
DocVQA (Mathew et al., 2021) LMMs-Eval Docs and Infographics $10.5$K ANLS Unknown
MMBench (Liu et al., 2023b) LMMs-Eval Multi-disciplinary $24$K GPT Apache-2.0
MMVET (Yu et al., 2024) LMMs-Eval Multi-disciplinary $218$ GPT Apache-2.0
MP-DocVQA (Tito et al., 2023) LMMs-Eval Docs and Infographics $5.2$K QEM MIT
NoCaps (Agrawal et al., 2019) LMMs-Eval Captioning $4.5$K ROUGE MIT
OK-VQA (Marino et al., 2019) LMMs-Eval VQA $5.1$K ANLS Unknown
RefCOCO (Kazemzadeh et al., 2014; Mao et al., 2016) LMMs-Eval Captioning $38$K ROUGE Apache-2.0
ScienceQA (Lu et al., 2022) LMMs-Eval Maths+Science $12.6$K EM CC BY-NC-SA 4.0
TextCaps (Sidorov et al., 2020) LMMs-Eval Captioning $3.2$K ROUGE CC BY 4.0
TextVQA (Singh et al., 2019) LMMs-Eval VQA $5$K EM CC BY 4.0
COCO (Lin et al., 2014) VHELM+LMMs-Eval Captioning $45.5$K ROUGE CC-BY-4.0
Flickr30k (Young et al., 2014) VHELM+LMMs-Eval Captioning $31$K ROUGE CC-0 Public Domain
GQA(Hudson and Manning, 2019) VHELM+LMMs-Eval Scene Understanding $12.6$K QEM CC-BY-4.0
MathVista (Lu et al., 2024a) VHELM+LMMs-Eval Maths+Science $1$K QEM/GPT4 CC-BY-SA-4.0
MME (Fu et al., 2023) VHELM+LMMs-Eval Multi-disciplinary $2.4$K QEM/C+P Unknown
MMMU (Yue et al., 2024) VHELM+LMMs-Eval Multi-disciplinary $900$ QEM CC BY-SA 4.0
POPE (Li et al., 2023b) VHELM+LMMs-Eval Hallucination $9$K QEM/EM MIT
SEED-Bench (Li et al., 2023a, 2024a) VHELM+LMMs-Eval Multi-disciplinary $42.5$K QEM/EM Apache
VizWiz (Gurari et al., 2018) VHELM+LMMs-Eval VQA $4.3$K QEM/EM CC BY 4.0
VQAv2 (Goyal et al., 2017) VHELM+LMMs-Eval VQA $214$K QEM/EM CC BY 4.0
@@
Vision Arena (Lu et al., 2024b) - Pairwise Battles $9$K - MIT
LMMs-Eval(Prometheus2) (Kim et al., 2024) - Pairwise Battles $610$K - MIT
@@
Table 6: Datasets in ONEBench-LMM: a diverse collection of benchmarks testing the abilities of LLMs in tasks such as general VQA, image captioning, hate speech detection, bias and hallucination understanding, maths and science, documents and infographics, scene understanding and sequential reasoning as well as the performance of LMMs in pairwise battles. Additional preference comparisons are sampled randomly from LMMs-Eval, which are excluded from the cardinal measurement sample pool.
Appendix B Models used in ONEBench: Further Details